Find the first negative term of the sequence 999,995,991,987
 
Please give me answer in full description
 

Grade:11

1 Answers

User
5 years ago
a=999, d=-4, tn=a+(n-1) d, let tn=-1, so 999+(n-1) (-4) =-1, (n-1) (-4) =-1-999, (n-1) =-1000/-4, so(n-1) =250, n=250+1=251, so n=251